OpenStudy (anonymous):

is this less or more 14.006 14.06

OpenStudy (anonymous):

14.06 is more than 14.006

ganeshie8 (ganeshie8):

yes 14.06 = 14060/1000 14.006 = 14006/1000

OpenStudy (anonymous):

Yes what Ganeshie8 said.

OpenStudy (anonymous):

14.06 is more than 14.006